OpenStudy (anonymous):

Algebra question just 1

OpenStudy (anonymous):

no its not letting me

OpenStudy (jhannybean):

\[\sqrt{3}+\sqrt{\frac{ 4 }{ 3}}\] ?

OpenStudy (anonymous):

i thought we had to simplify \[\sqrt{3+4} and \sqrt{3}?\]

OpenStudy (jhannybean):